# SQuat: Subspace-orthogonal KV Cache Quantization

### v1

@@ Abstract
The key-value (KV) cache accelerates LLMs decoding by storing KV tensors from previously generated tokens. It reduces redundant computation at the cost of increased memory usage. To mitigate this overhead, existing approaches compress KV tensors into lower-bit representations; however, quantization errors can accumulate as more tokens are generated, potentially resulting in undesired outputs. In this paper, we introduce SQuat (Subspace-orthogonal KV cache quantization). It first constructs a subspace spanned by query tensors to capture the most critical task-related information. During key tensor quantization, it enforces that the difference between the (de)quantized and original keys remains orthogonal to this subspace, minimizing the impact of quantization errors on the attention mechanism’s outputs. SQuat requires no model fine-tuning, no additional calibration dataset for offline learning, and is grounded in a theoretical framework we develop. Through numerical experiments, we show that our method reduces peak memory by $\mathbf{2.17\times}\sim\mathbf{2.82\times}$, improves throughput by $\mathbf{2.45\times}\sim\mathbf{3.60\times}$, and achieves more favorable benchmark scores than existing KV cache quantization algorithms.
@@ 1.1 Related Work
Finally, we defer a discussion of broader related work to Appendix A.
@@ 2 Background
For the sake of illustration, we assume batch size is $1$ and the attention mechanism has a single head. A transformer consists of $L$ layers and each layer contains a multi-head attention and a feed-forward network. Suppose the input token embeddings are: $\mathbf{X}\in\mathbb{R}^{n\times d}$.
@@
In the prefill phase, let $\mathbf{X}_{0}\in\mathbb{R}^{l_{\text{prompt}}\times d}$ represent the input tensor, where $l_{\text{prompt}}$ is the length of the user’s input prompt and $d$ denotes the model’s hidden size. During this phase, the key ($\mathbf{K}$) and value ($\mathbf{V}$) tensors are computed and cached. This is followed by the decoding phase, in which the LLM predicts the next token sequentially, conditioned on the tokens generated so far. Note that LLMs only compute the query, key, and value for the newly generated token. The new key and value are then concatenated with the previously cached tensors. This caching mechanism significantly reduces computational overhead, as it avoids recomputing the key and value tensors for all prior tokens at every decoding step, resulting in faster inference. However, the KV cache may become a memory bottleneck when handling long contexts or large batch sizes (see Appendix C). Compressing the KV cache—through methods such as quantization—offers a natural solution to this challenge.
@@ 3.2 Exploring Query Subspace
For a query tensor $\bm{q}$, we define its deviation from a given subspace $\mathcal{P}$ as $\|\bm{q}-\mathsf{Proj}_{\mathcal{P}}(\bm{q})\|_{2}$, where $\mathsf{Proj}_{\mathcal{P}}(\bm{q})$ denotes the projection of $\bm{q}$ onto $\mathcal{P}$. When the query tensor is normalized, its deviation from any subspace stays within $[0,1]$.
@@
Figure 2 (left) illustrates the deviation of (normalized) query tensors of tokens in the math-500 sequence from the four subspaces. As shown, a 30-dimensional subspace can capture the entire sequence’s query tensors with ~$20\%$ deviation error. Moreover, using prompt-only query tensors—available before the decoding phase begins—or query tensors from a different math-500 sequence does not increase the deviation error. The deviation increases slightly, however, when the subspace is spanned using a sequence from QuALITY, a different dataset. Based on these observations, we apply SVD to the query tensors of all prompt tokens. Then we define the task-relevant query subspace as $\hat{\mathbf{Q}}\in\mathbb{R}^{r\times d}$, constructed from the top $r$ singular vectors scaled by their corresponding singular values.
Figure 2: Left: query tensors tend to lie in a low-dimensional subspace. This subspace can be identified using query tensors from prompt tokens, which are available before decoding begins. Middle and Right: the absolute difference in attention scores between using original and quantized key tensors. As shown, our method better preserves attention scores. See Appendix E.1 for experimental setup.
@@ 3.3 Key Tensors Quantization
We propose an iterative algorithm to solve (1). In each iteration, we selects a coordinate [^1] In Appendix B.2, we extend our algorithm to do block-by-block quantization of key tensors, rather than element-by-element. In our experiments, each key tensor is partitioned into no more than 4 blocks, leading to at most 4 iterations. of $\bm{k}$, quantize it, and update the remaining coordinates to ensure the residual is orthogonal to the query subspace as much as possible. The process repeats for the remaining coordinates until all coordinates of $\bm{k}$ are quantized.
@@ 4 Numerical Experiments
We consider FP16 (no quantization) and existing KV cache quantization algorithms—KIVI (Liu et al., 2024e), GEAR (Kang et al., 2024), ZipCache (He et al., 2024)—which, like our method, do not require fine-tuning or a calibration dataset. We evaluate these methods on four open-source LLMs: Llama-2-7B, Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, DeepSeek-R1-Distill-Llama-8B. We assess performance across several reasoning and long-context benchmarks, and report the KV size as the average percentage of the compressed cache relative to the FP16 cache at the end of generation.
@@
We evaluate SQuat and baseline methods on long-context benchmarks, following the recommendations of Liu et al. (2024e) and using the same set of tasks (Qasper, QMSum, MultiNews, TREC, TriviaQA, SAMSum, LCC, and RepoBench-P) from LongBench (Bai et al., 2023). These tasks evaluate the ability of LLMs to process and reason over long contexts, such as document-level question answering and summarization. As shown in Table 2, SQuat achieves higher average scores across all models than other quantization baselines and are nearly lossless compared with FP16.
@@
We apply SQuat to quantize the KV cache to 2-bit and evaluate its efficiency in terms of memory usage, latency, and throughput on the Llama-2-7B model. Empirically, GEAR and ZipCache run much slower than KIVI (Liu et al., 2024e) and SQuat, so we compare SQuat only against KIVI and an FP16 baseline (without quantization) in Figure 3. To ensure a fair comparison, we adopt the same experimental setup as KIVI, which synthesizes workloads based on ShareGPT (2023). We evaluate two residual buffer lengths (i.e., the number of most recent tokens whose KV cache is stored in full precision), 32 and 128, for both our algorithm and KIVI. We gradually increase the batch size until the GPU runs out of memory. All experiments are conducted on a single Nvidia H100 GPU (80GB). As shown in Figure 3, SQuat requires the same peak GPU memory as KIVI, and reduces peak memory by $\mathbf{2.17\times\sim 2.82\times}$ compared to FP16. SQuat enables up to a 4×4$\times$4 × increase in batch size compared to the FP16 baseline, leading to a throughput improvement of $\mathbf{2.45\times\sim 3.60\times}$, comparable to KIVI’s throughput improvement of $\mathbf{2.63\times\sim 3.82\times}$.
More experiments.
We provide an overview of additional results in appendices. First, we present the experimental setup, including the choice of hyper-parameters for SQuat, SQuatpre, and other baseline methods (Appendix E.2). Second, we present an ablation study that suggests hyperparameter selection for SQuat (Appendix E.3). Finally, we present evaluation results of DeepSeek-R1-Distill-Llama-8B on the math-500 dataset (Appendix E.4). This experiment highlights a key use case for KV cache quantization: modern reasoning models often require long responses to “think” through problems—e.g., by generating chain-of-thought reasoning—before arriving at a final answer. In such scenarios, the KV cache becomes a memory bottleneck, and quantization proves especially helpful when compute resources are limited.
@@ 5 Conclusion, Future work, and Limitations
Appendix A More Related Work
Efficient system and architecture designs.
Model compression.
Appendix B Omitted Proofs
B.1 Proof of Theorem 1
@@
Note that $\mathsf{softmax}$ is $1/2$-Lipschitz continuous (see Appendix A.4 in Alghamdi et al., 2022). Therefore,
@@
B.2 Block-Wise Quantization
@@
Now, we consider quantizing a block of $g$ elements per iteration in a sequential manner. Suppose the hidden dimension $d$ of the key tensors is divisible by $g$. At iteration $t$, the first $(t-1)g$ elements have already been quantized and are kept fixed. During this step, we quantize the elements indexed from $(t-1)g+1$ to $tg$ and update the rest of the elements by using the following lemma.
@@
where $\bm{d}\in\mathbb{R}^{g}$ has its $i$-th element corresponding to the $(t-1)g+i$-th element of $\mathsf{deq}(\mathsf{qtz}(\hat{\bm{k}}_{t-1}))-\hat{\bm{k}}_{t-1}$.
@@
We denote $\boldsymbol{\delta}=\hat{\bm{k}}_{t}-\hat{\bm{k}}_{t-1}$ and let $\bm{d}\in\mathbb{R}^{g}$ has its $i$-th element corresponding to the $(t-1)g+i$-th element of $\mathsf{deq}(\mathsf{qtz}(\hat{\bm{k}}_{t-1}))-\hat{\bm{k}}_{t-1}$. Moreover, we denote
@@
Appendix C Memory Requirements for KV Cache Management
@@
Appendix D Details of Our Main Algorithm
@@
We assume a batch size of $1$ in the pseudocode. When the batch size exceeds $1$, we share the same matrices $\mathbf{A}_{1}^{-1},\dots,\mathbf{A}_{T-1}^{-1},\mathbf{P}_{\text{inv}}$ across all samples in the batch, computing them from the first sample. This significantly reduces memory usage for large batch sizes. Empirically, we find that this does not degrade LLM performance compared to computing separate matrices for each sample. This aligns with our observation in Section 3.2 that query tensors tend to lie within a small subspace, even when constructed from different sequences in the same dataset.
@@
5:$\mathbf{Q}=\mathbf{X}\mathbf{W}^{Q}$, $\mathbf{K}=\mathbf{X}\mathbf{W}^{K}$
6:_$\_,\boldsymbol{\Sigma},\mathbf{V}=\mathsf{SVD}(\mathbf{Q})$triangleright can be accelerated using block power method (Bentbib & Kanber, 2015)
@@
33:function Quant($\mathbf{K}$, $\mathbf{A}_{1}^{-1},\cdots,\mathbf{A}_{T-1}^{-1}$, $\mathbf{P}_{\text{inv}}$)
@@
Appendix E More Experimental Results
E.1 Experimental Setup for Figure 2
@@
E.2 Experimental Setup for Section 4
For SQuat, SQuatpre, we set the group size (i.e., the number of elements quantized together) to 32 and the residual length (i.e., the number of most recent tokens whose KV cache is stored in full precision) to 32. The KV cache is quantized to 2 bits. This configuration provides the best throughput and minimal memory usage (see Figure 3). For LM-Eval tasks and all LLMs, we set the subspace dimension to $r=5$, the regularization coefficient to $\lambda=0.001$, and the number of elements to quantize per iteration to $g=64$, resulting in just two iterations of our algorithm. For LongBench tasks, which involve longer inputs, we adjust the hyper-parameters as follows. For Llama-2-7B-hf, we use $r=5$, $\lambda=0.001$, $g=32$ for SQuatpre, and $r=40$, $\lambda=0.001$, $g=64$ for SQuat. For Llama-3.1-8B-Instruct, we set SQuatpre to $r=60$, $\lambda=0.0005$, $g=32$, and SQuat to $r=10$, $\lambda=0.0005$, $g=64$. For Mistral-7B-Instruct-v0.3, we use $r=20$, $\lambda=0.0005$, $g=64$ for SQuatpre, and $r=5$, $\lambda=0.001$, $g=64$ for SQuat.
For KIVI and GEAR, we use the same configuration: a group size of 32, a residual length of 32, and 2-bit quantization for the KV cache. For KIVI, we use the official implementation from their GitHub repository. For GEAR, we set rank = rankv = 2 and loop = 3, following the recommendations in their test script. Since their CUDA-based implementation does not support the Mistral model, we compare only on LLaMA models using their implementation from GitHub repository. For ZipCache, we quantize important tokens to 4 bits and unimportant tokens to 2 bits, with streaming_gap = 32 to match the setup used for other baselines and our method. We set the unimportant_ratio to 0.6 for Llama-2-7B on LM-eval tasks, 0.7 on LongBench tasks, 0.5 for Llama-3.1-8B-Instruct on LM-eval tasks, and 0.6 on LongBench tasks. These values ensure that their KV cache size is roughly aligned with that of other baselines, allowing for a fair comparison. As ZipCache is only implemented for LLaMA in their GitHub repository, we include it in comparisons for LLaMA models only.
@@ E.3 Ablation Study
Figure 6: Comparison of different KV cache quantization methods and FP16 (no quantization) on the math-500 dataset using the DeepSeek-R1-Distill-Llama-8B model.
@@
E.4 Results on math-500
We evaluate DeepSeek-R1-Distill-Llama-8B on the math-500 dataset (Lightman et al., 2023), setting the maximum number of new tokens to 8k to ensure the model has sufficient window for multi-stage reasoning. This setup highlights a key use case for KV cache quantization: modern reasoning models often require long responses to “think” through problems—e.g., by generating chain-of-thought reasoning—before arriving at a final answer. In such scenarios, the KV cache becomes a memory bottleneck, and quantization proves especially helpful when compute resources are limited. As shown in Figure 6, SQuatpre outperforms all existing KV cache quantization methods.

### v2

@@ Abstract
The key-value (KV) cache accelerates LLMs decoding by storing KV tensors from previously generated tokens. It reduces redundant computation at the cost of increased memory usage. To mitigate this overhead, existing approaches compress KV tensors into lower-bit representations; however, quantization errors can accumulate as more tokens are generated, potentially resulting in undesired outputs. In this paper, we introduce SQuat (Subspace-orthogonal KV cache quantization). It first constructs a subspace spanned by query tensors to capture the most critical task-related information. During key tensor quantization, it enforces that the difference between the (de)quantized and original keys remains orthogonal to this subspace, minimizing the impact of quantization errors on the attention mechanism’s outputs. SQuat requires no model fine-tuning, no additional calibration dataset for offline learning, and is grounded in a theoretical framework we develop. Through numerical experiments, we show that our method reduces peak memory by $\mathbf{2.17\times}\sim\mathbf{2.82\times}$, improves throughput by $\mathbf{2.45\times}\sim\mathbf{3.60\times}$, and achieves more favorable benchmark scores than existing KV cache quantization algorithms. Our code is available at: https://github.com/Red-Hat-AI-Innovation-Team/SQuat
@@ 1.1 Related Work
Efficient system and architecture designs.
Model compression.
@@ 2 Background
For the sake of illustration, we assume batch size is 111 and the attention mechanism has a single head. A transformer consists of $L$ layers and each layer contains a multi-head attention and a feed-forward network. Suppose the input token embeddings are: $\mathbf{X}\in\mathbb{R}^{n\times d}$.
@@
In the prefill phase, let $\mathbf{X}_{0}\in\mathbb{R}^{l_{\text{prompt}}\times d}$ represent the input tensor, where $l_{\text{prompt}}$ is the length of the user’s input prompt and $d$ denotes the model’s hidden size. During this phase, the key ($\mathbf{K}$) and value ($\mathbf{V}$) tensors are computed and cached. This is followed by the decoding phase, in which the LLM predicts the next token sequentially, conditioned on the tokens generated so far. Note that LLMs only compute the query, key, and value for the newly generated token. The new key and value are then concatenated with the previously cached tensors. This caching mechanism significantly reduces computational overhead, as it avoids recomputing the key and value tensors for all prior tokens at every decoding step, resulting in faster inference. However, the KV cache may become a memory bottleneck when handling long contexts or large batch sizes (see Appendix B). Compressing the KV cache—through methods such as quantization—offers a natural solution to this challenge.
@@ 3.2 Exploring Query Subspace
For a query tensor $\bm{q}$, we define its deviation from a given subspace $\mathcal{P}$ as $\|\bm{q}-\mathsf{Proj}_{\mathcal{P}}(\bm{q})\|_{2}$, where $\mathsf{Proj}_{\mathcal{P}}(\bm{q})$ denotes the projection of $\bm{q}$ onto $\mathcal{P}$. When the query tensor is normalized, its deviation from any subspace stays within [0,1][0,1][ 0 , 1 ].
@@
[…]
Figure 2: Left: query tensors tend to lie in a low-dimensional subspace. This subspace can be identified using query tensors from prompt tokens, which are available before decoding begins. Middle and Right: the absolute difference in attention scores between using original and quantized key tensors. As shown, our method better preserves attention scores. See Appendix D.1 for experimental setup.
@@ 3.3 Key Tensors Quantization
We propose an iterative algorithm to solve (1). In each iteration, we selects a coordinate [^1] In Appendix A.2, we extend our algorithm to do block-by-block quantization of key tensors, rather than element-by-element. In our experiments, each key tensor is partitioned into no more than 4 blocks, leading to at most 4 iterations. of $\bm{k}$, quantize it, and update the remaining coordinates to ensure the residual is orthogonal to the query subspace as much as possible. The process repeats for the remaining coordinates until all coordinates of $\bm{k}$ are quantized.
@@ 4 Numerical Experiments
We consider FP16 (no quantization) and existing KV cache quantization algorithms—KIVI (Liu et al., 2024e), GEAR (Kang et al., 2024), ZipCache (He et al., 2024)—which, like our method, do not require fine-tuning or a calibration dataset. We evaluate these methods on four open-source LLMs: Llama-2-7B, Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, DeepSeek-R1-Distill-Llama-8B. We assess performance across several reasoning and long-context benchmarks, and report the KV size as the average percentage of the compressed cache relative to the FP16 cache at the end of generation. We present the experimental setup, including the choice of hyper-parameters for SQuat, SQuatpre, and other baseline methods in Appendix D.2.
@@
We evaluate SQuat and baseline methods on long-context benchmarks, following the recommendations of Liu et al. (2024e) and using the same set of tasks (Qasper, QMSum, MultiNews, TREC, TriviaQA, SAMSum, LCC, and RepoBench-P) from LongBench (Bai et al., 2023). These tasks evaluate the ability of LLMs to process and reason over long contexts, such as document-level question answering and summarization, and require deep understanding across diverse real-world scenarios. As shown in Table 2, SQuat achieves higher average scores across all models than other quantization baselines and are nearly lossless compared with FP16.
@@
Reasoning model.
Figure 3: Experiments using a reasoning model (DeepSeek-R1-Distill-Llama-8B) on the math-500 dataset. We compare SQuat and SQuatpre with different KV cache quantization baselines and FP16 (no quantization).
We evaluate DeepSeek-R1-Distill-Llama-8B on the math-500 dataset (Lightman et al., 2023). We set the maximum number of new tokens to 8k to ensure the model has sufficient window for multi-stage reasoning. This setup highlights a key use case for KV cache quantization: modern reasoning models often require long responses to “think” through problems—e.g., by generating chain-of-thought reasoning—before arriving at a final answer. In such scenarios, the KV cache becomes a memory bottleneck, and quantization proves especially helpful when compute resources are limited. As shown in Figure 3, both SQuat and SQuatpre outperform all existing KV cache quantization baselines.
@@
We apply SQuat to quantize the KV cache to 2-bit and evaluate its efficiency in terms of memory usage, latency, and throughput on the Llama-2-7B model. Empirically, GEAR and ZipCache run much slower than KIVI (Liu et al., 2024e) and SQuat, so we compare SQuat only against KIVI and an FP16 baseline (without quantization) in Figure 4. To ensure a fair comparison, we adopt the same experimental setup as KIVI, which synthesizes workloads based on ShareGPT (2023). We evaluate two residual buffer lengths (i.e., the number of most recent tokens whose KV cache is stored in full precision), 32 and 128, for both our algorithm and KIVI. We gradually increase the batch size until the GPU runs out of memory. All experiments are conducted on a single Nvidia H100 GPU (80GB). As shown in Figure 4, SQuat reduces peak GPU memory by $\mathbf{2.17\times}\sim\mathbf{2.82\times}$. This memory efficiency enables up to a 4×4$\times$4 × increase in batch size compared to the FP16 baseline, leading to a throughput improvement of $\mathbf{2.45\times\sim 3.60\times}$, which is comparable to KIVI’s throughput improvement of $\mathbf{2.63\times\sim 3.82\times}$.
@@ 5 Conclusion, Future work, and Limitations
Acknowledgment
We would like to thank Shivchander Sudalairaj, Abhishek Bhandwaldar, Mustafa Eyceoz, Aldo Pareja from the Red Hat AI Innovation Team for their help in setting up some of the evaluation benchmarks.
Appendix A Omitted Proofs
A.1 Proof of Theorem 1
@@
Note that $\mathsf{softmax}$ is 1/21/21 / 2-Lipschitz continuous (see Appendix A.4 in Alghamdi et al., 2022). Therefore,
@@
A.2 Block-Wise Quantization
@@
[…]
@@
[…]
@@
[…]
@@
Appendix B Memory Requirements for KV Cache Management
@@
Appendix C Details of Our Main Algorithm
@@
We assume a batch size of 111 in the pseudocode. When the batch size exceeds 111, we share the same matrices $\mathbf{A}_{1}^{-1},\dots,\mathbf{A}_{T-1}^{-1},\mathbf{P}_{\text{inv}}$ across all samples in the batch, computing them from the first sample. This significantly reduces memory usage for large batch sizes. Empirically, we find that this does not degrade LLM performance compared to computing separate matrices for each sample. This aligns with our observation in Section 3.2 that query tensors tend to lie within a small subspace, even when constructed from different sequences in the same dataset.
@@
[…]
@@
Appendix D More Experimental Results
D.1 Experimental Setup for Figure 2
@@
D.2 Experimental Setup for Section 4
For SQuat, SQuatpre, we set the group size (i.e., the number of elements quantized together) to 32 and the residual length (i.e., the number of most recent tokens whose KV cache is stored in full precision) to 32. The KV cache is quantized to 2 bits. This configuration provides the best throughput and minimal memory usage (see Figure 4). For LM-Eval tasks and all LLMs, we set the subspace dimension to $r=5$, the regularization coefficient to $\lambda=0.001$, and the number of elements to quantize per iteration to $g=64$, resulting in just two iterations of our algorithm. For LongBench tasks, which involve longer inputs, we adjust the hyper-parameters as follows. For Llama-2-7B-hf, we use $r=5$, $\lambda=0.001$, $g=32$ for SQuatpre, and $r=40$, $\lambda=0.001$, $g=64$ for SQuat. For Llama-3.1-8B-Instruct, we set SQuatpre to $r=60$, $\lambda=0.0005$, $g=32$, and SQuat to $r=10$, $\lambda=0.0005$, $g=64$. For Mistral-7B-Instruct-v0.3, we use $r=20$, $\lambda=0.0005$, $g=64$ for SQuatpre, and $r=5$, $\lambda=0.001$, $g=64$ for SQuat. For our experiments using DeepSeek-R1-Distill-Llama-8B on the math-500 dataset (Figure 3), we set $r=40$, $\lambda=0.0005$, and $g=64$ for both SQuat and SQuatpre.
For KIVI and GEAR, we use the same configuration: a group size of 32, a residual length of 32, and 2-bit quantization for the KV cache. For KIVI, we use the official implementation from their GitHub repository. For GEAR, we set rank = rankv = 2 and loop = 3, following the recommendations in their test script. Since their CUDA-based implementation does not support the Mistral model, we compare only on LLaMA models using their implementation from GitHub repository. For ZipCache, we quantize important tokens to 4 bits and unimportant tokens to 2 bits, with streaming_gap = 32 to match the setup used for other baselines and our method. We set the unimportant_ratio to 0.6 for Llama-2-7B on LM-eval tasks, 0.7 on LongBench tasks, 0.5 for Llama-3.1-8B-Instruct on LM-eval tasks, and 0.6 on LongBench tasks. These values ensure that their KV cache size is roughly aligned with that of other baselines, allowing for a fair comparison. As ZipCache is only implemented for LLaMA in their GitHub repository (as of the date of our paper submission), we include it in comparisons for LLaMA models only.